# When and How: Learning Identifiable Latent States for Nonstationary Time Series Forecasting

### v1

@@ Appendix D Evident Lower Bound
$\small\begin{split}&\ln p({\mathbf{x}}_{t+1:T},{\mathbf{x}}_{1:t})=\ln\frac{p({\mathbf{x}}_{t+1:T},{\mathbf{z}}^{e}_{1:T},{\mathbf{z}}^{s}_{1:T},{\mathbf{x}}_{1:t})}{p({\mathbf{z}}^{e}_{1:T},{\mathbf{z}}^{s}_{1:T}|{\mathbf{x}}_{1:t},{\mathbf{x}}_{t+1:T})}=\ln\frac{p({\mathbf{x}}_{t+1:T},{\mathbf{z}}^{e}_{1:t},{\mathbf{z}}^{s}_{1:t},{\mathbf{z}}^{e}_{t+1:T},{\mathbf{z}}^{s}_{t+1:T},{\mathbf{x}}_{1:t})}{p({\mathbf{z}}^{e}_{1:T},{\mathbf{z}}^{s}_{1:T}|{\mathbf{x}}_{1:T})}\\ \\ =&\mathbb{E}_{q({\mathbf{z}}^{s}_{1:t}|{\mathbf{x}}_{1:t})}\mathbb{E}_{q({\mathbf{z}}^{s}_{t+1:T}|{\mathbf{z}}^{s}_{1:t})}\mathbb{E}_{q({\mathbf{z}}^{e}_{1:t}|{\mathbf{x}}_{1:t})}\mathbb{E}_{q({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{z}}^{e}_{1:t})}\ln\frac{p({\mathbf{x}}_{t+1:T}|{\mathbf{z}}_{t+1:T}^{e},{\mathbf{z}}_{t+1:T}^{s})p({\mathbf{x}}_{1:t}|{\mathbf{z}}_{1:t}^{e},{\mathbf{z}}_{1:t}^{s})p({\mathbf{z}}^{s}_{1:t})p({\mathbf{z}}_{1:t}^{e})p({\mathbf{z}}^{s}_{t+1:T}|{\mathbf{z}}^{s}_{1:t})p({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{z}}^{e}_{1:t})}{q({\mathbf{z}}^{e}_{1:t}|{\mathbf{x}}_{1:t})q({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{z}}^{e}_{1:t})q({\mathbf{z}}^{s}_{1:t}|{\mathbf{x}}_{1:t})q({\mathbf{z}}_{t+1:T}^{s}|{\mathbf{z}}^{s}_{1:t})}\\ &+D_{KL}(q({\mathbf{z}}^{s}_{t+1:T}|{\mathbf{z}}^{s}_{1:t})||p({\mathbf{z}}^{s}_{t+1:T}|{\mathbf{z}}^{s}_{1:t},{\mathbf{x}}_{1:T},{\mathbf{z}}^{e}_{1:T}))+D_{KL}(q({\mathbf{z}}^{s}_{1:t}|{\mathbf{x}}_{1:t})||p({\mathbf{z}}^{s}_{1:t}|{\mathbf{x}}_{1:T},{\mathbf{z}}^{e}_{1:T}))\\ &+D_{KL}(q({\mathbf{z}}^{e}_{1:t}|{\mathbf{x}}_{1:t})||p({\mathbf{z}}^{e}_{1:t}|{\mathbf{x}}_{1:t}))+D_{KL}(q({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{z}}^{e}_{1:t})||p({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{x}}_{1:T},{\mathbf{z}}^{e}_{1:t}))\\ \\ \geq&\mathbb{E}_{q({\mathbf{z}}^{s}_{1:t}|{\mathbf{x}}_{1:t})}\mathbb{E}_{q({\mathbf{z}}^{s}_{t+1:T}|{\mathbf{z}}^{s}_{1:t})}\mathbb{E}_{q({\mathbf{z}}^{e}_{1:t}|{\mathbf{x}}_{1:t})}\mathbb{E}_{q({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{z}}^{e}_{1:t})}\ln\frac{p({\mathbf{x}}_{t+1:T}|{\mathbf{z}}_{t+1:T}^{e},{\mathbf{z}}_{t+1:T}^{s})p({\mathbf{z}}^{s}_{1:t}|{\mathbf{x}}_{1:t},{\mathbf{z}}^{e}_{1:t})p({\mathbf{z}}^{s}_{t+1:T}|{\mathbf{z}}^{s}_{1:t})p({\mathbf{z}}^{e}_{1:t}|{\mathbf{x}}_{1:t})p({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{z}}^{e}_{1:t})}{q({\mathbf{z}}^{e}_{1:t}|{\mathbf{x}}_{1:t})q({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{z}}^{e}_{1:t})q({\mathbf{z}}^{s}_{1:t}|{\mathbf{x}}_{1:t})q({\mathbf{z}}_{t+1:T}^{s}|{\mathbf{z}}^{s}_{1:t})}\\ \\ =&\underbrace{\mathbb{E}_{q({\mathbf{z}}^{s}_{1:t}|{\mathbf{x}}_{1:t})}\mathbb{E}_{q({\mathbf{z}}^{e}_{1:t}|{\mathbf{x}}_{1:t})}\ln p({\mathbf{x}}_{1:t}|{\mathbf{z}}^{s}_{1:t},{\mathbf{z}}^{e}_{1:t})}_{\mathcal{L}_{rec}}+\underbrace{\mathbb{E}_{q({\mathbf{z}}^{s}_{1:t}|{\mathbf{x}}_{1:t})}\mathbb{E}_{q({\mathbf{z}}^{s}_{t+1:T}|{\mathbf{z}}^{s}_{1:t})}\mathbb{E}_{q({\mathbf{z}}^{e}_{1:t}|{\mathbf{x}}_{1:t})}\mathbb{E}_{q({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{z}}^{e}_{1:t})}\ln p({\mathbf{x}}_{t+1:T}|{\mathbf{z}}^{s}_{t+1:T},{\mathbf{z}}^{e}_{t+1:T})}_{\mathcal{L}_{pre}}\\ &-\underbrace{D_{KL}(q({\mathbf{z}}^{s}_{1:t}|{\mathbf{x}}_{1:t})||p({\mathbf{z}}^{s}_{1:t}))-\mathbb{E}_{q({\mathbf{z}}^{s}_{1:t}|{\mathbf{x}}_{1:t})}\Big{[}D_{KL}(q({\mathbf{z}}^{s}_{t:1+T}|{\mathbf{z}}^{s}_{1:t})||p({\mathbf{z}}^{s}_{t+1:T}|{\mathbf{z}}^{s}_{1:t}))\Big{]}}_{\mathcal{L}_{KLD}^{s}}\\ &-\underbrace{D_{KL}(q({\mathbf{z}}^{e}_{1:t}|{\mathbf{x}}_{1:t})||p({\mathbf{z}}^{e}_{1:t}))-\mathbb{E}_{q({\mathbf{z}}^{e}_{1:t}|{\mathbf{x}}_{1:t})}\Big{[}D_{KL}(q({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{z}}^{e}_{1:t})||p({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{z}}^{e}_{1:t}))\Big{]}}_{\mathcal{L}_{KLD}^{e}}\end{split}$ (24)
@@ E.1 Identification of Latent Domain Variables ${\mathbf{u}}_{t}$
(Theorem 9 in (Allman et al., 2009)) Let $\mathbb{P}$ be a mixture in the form of Equation (26), such that for every $j$, the measures $\mu_{i,j}$ are linearly independent. Then, if $c\geq 3$, $\{\pi_{i},\mu_{i,j}\}$ are identifiable from $\mathbb{P}$ up to label swapping.
@@
The proof of this lemma can refer to Theorem 9 of (Allman et al., 2009). In general, Lemma E.1 shows that if the joint distribution of observation $\mathbb{P}$ can be decomposed into three linearly independent measures w.r.t. $\mu_{i,j}$ as shown in Equation (26), then the distributions of discrete latent variables are identifiable. Based on this Lemma, we further show the identification results of latent environments as follows.
@@
(Identification of the latent environment ${\mathbf{e}}_{t}$.) Suppose the observed data is generated following the data generation process in Figure 2 and Equation (1)-(3). Then we further make the following assumptions:
@@
  • A3.1.1 (Prior Environment Number:) The number of latent environments of the Markov process, $E$, is known.
  • A3.1.2 (Full Rank:) The transition matrix $\mathbf{A}$ is full rank.
@@
  • A3.1.3 (Linear Independence:) For $e=1,2,\cdots,E$, the probability measures $\mu_{e}=p({\mathbf{x}}_{t}|{\mathbf{z}}_{t}^{e},e)$ are linearly independence and for any two different probability measures $\mu_{i},\mu_{j}$, their ratio $\frac{\mu_{i}}{\mu_{j}}$ are linearly independence.
  • A3.1.4 (Sufficient Observations:) For each ${\mathbf{e}}=\{1,\cdots,E\}$, there are at least 2 consecutive observations.

### v4

@@ Appendix D Evident Lower Bound
$\tiny\begin{split}&\ln p({\mathbf{x}}_{t+1:T},{\mathbf{x}}_{1:t})=\ln\frac{p({\mathbf{x}}_{t+1:T},{\mathbf{z}}^{e}_{1:T},{\mathbf{z}}^{s}_{1:T},{\mathbf{x}}_{1:t})}{p({\mathbf{z}}^{e}_{1:T},{\mathbf{z}}^{s}_{1:T}|{\mathbf{x}}_{1:t},{\mathbf{x}}_{t+1:T})}=\ln\frac{p({\mathbf{x}}_{t+1:T},{\mathbf{z}}^{e}_{1:t},{\mathbf{z}}^{s}_{1:t},{\mathbf{z}}^{e}_{t+1:T},{\mathbf{z}}^{s}_{t+1:T},{\mathbf{x}}_{1:t})}{p({\mathbf{z}}^{e}_{1:T},{\mathbf{z}}^{s}_{1:T}|{\mathbf{x}}_{1:T})}\\ \\ =&\mathbb{E}_{q({\mathbf{z}}^{s}_{1:t}|{\mathbf{x}}_{1:t})}\mathbb{E}_{q({\mathbf{z}}^{s}_{t+1:T}|{\mathbf{z}}^{s}_{1:t})}\mathbb{E}_{q({\mathbf{z}}^{e}_{1:t}|{\mathbf{x}}_{1:t})}\mathbb{E}_{q({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{z}}^{e}_{1:t})}\ln\frac{p({\mathbf{x}}_{t+1:T}|{\mathbf{z}}_{t+1:T}^{e},{\mathbf{z}}_{t+1:T}^{s})p({\mathbf{x}}_{1:t}|{\mathbf{z}}_{1:t}^{e},{\mathbf{z}}_{1:t}^{s})p({\mathbf{z}}^{s}_{1:t})p({\mathbf{z}}_{1:t}^{e})p({\mathbf{z}}^{s}_{t+1:T}|{\mathbf{z}}^{s}_{1:t})p({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{z}}^{e}_{1:t})}{q({\mathbf{z}}^{e}_{1:t}|{\mathbf{x}}_{1:t})q({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{z}}^{e}_{1:t})q({\mathbf{z}}^{s}_{1:t}|{\mathbf{x}}_{1:t})q({\mathbf{z}}_{t+1:T}^{s}|{\mathbf{z}}^{s}_{1:t})}\\ &+D_{KL}(q({\mathbf{z}}^{s}_{t+1:T}|{\mathbf{z}}^{s}_{1:t})||p({\mathbf{z}}^{s}_{t+1:T}|{\mathbf{z}}^{s}_{1:t},{\mathbf{x}}_{1:T},{\mathbf{z}}^{e}_{1:T}))+D_{KL}(q({\mathbf{z}}^{s}_{1:t}|{\mathbf{x}}_{1:t})||p({\mathbf{z}}^{s}_{1:t}|{\mathbf{x}}_{1:T},{\mathbf{z}}^{e}_{1:T}))\\ &+D_{KL}(q({\mathbf{z}}^{e}_{1:t}|{\mathbf{x}}_{1:t})||p({\mathbf{z}}^{e}_{1:t}|{\mathbf{x}}_{1:t}))+D_{KL}(q({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{z}}^{e}_{1:t})||p({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{x}}_{1:T},{\mathbf{z}}^{e}_{1:t}))\\ \\ \geq&\mathbb{E}_{q({\mathbf{z}}^{s}_{1:t}|{\mathbf{x}}_{1:t})}\mathbb{E}_{q({\mathbf{z}}^{s}_{t+1:T}|{\mathbf{z}}^{s}_{1:t})}\mathbb{E}_{q({\mathbf{z}}^{e}_{1:t}|{\mathbf{x}}_{1:t})}\mathbb{E}_{q({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{z}}^{e}_{1:t})}\ln\frac{p({\mathbf{x}}_{t+1:T}|{\mathbf{z}}_{t+1:T}^{e},{\mathbf{z}}_{t+1:T}^{s})p({\mathbf{z}}^{s}_{1:t}|{\mathbf{x}}_{1:t},{\mathbf{z}}^{e}_{1:t})p({\mathbf{z}}^{s}_{t+1:T}|{\mathbf{z}}^{s}_{1:t})p({\mathbf{z}}^{e}_{1:t}|{\mathbf{x}}_{1:t})p({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{z}}^{e}_{1:t})}{q({\mathbf{z}}^{e}_{1:t}|{\mathbf{x}}_{1:t})q({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{z}}^{e}_{1:t})q({\mathbf{z}}^{s}_{1:t}|{\mathbf{x}}_{1:t})q({\mathbf{z}}_{t+1:T}^{s}|{\mathbf{z}}^{s}_{1:t})}\\ \\ =&\underbrace{\mathbb{E}_{q({\mathbf{z}}^{s}_{1:t}|{\mathbf{x}}_{1:t})}\mathbb{E}_{q({\mathbf{z}}^{e}_{1:t}|{\mathbf{x}}_{1:t})}\ln p({\mathbf{x}}_{1:t}|{\mathbf{z}}^{s}_{1:t},{\mathbf{z}}^{e}_{1:t})}_{\mathcal{L}_{rec}}+\underbrace{\mathbb{E}_{q({\mathbf{z}}^{s}_{1:t}|{\mathbf{x}}_{1:t})}\mathbb{E}_{q({\mathbf{z}}^{s}_{t+1:T}|{\mathbf{z}}^{s}_{1:t})}\mathbb{E}_{q({\mathbf{z}}^{e}_{1:t}|{\mathbf{x}}_{1:t})}\mathbb{E}_{q({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{z}}^{e}_{1:t})}\ln p({\mathbf{x}}_{t+1:T}|{\mathbf{z}}^{s}_{t+1:T},{\mathbf{z}}^{e}_{t+1:T})}_{\mathcal{L}_{pre}}\\ &-\underbrace{D_{KL}(q({\mathbf{z}}^{s}_{1:t}|{\mathbf{x}}_{1:t})||p({\mathbf{z}}^{s}_{1:t}))-\mathbb{E}_{q({\mathbf{z}}^{s}_{1:t}|{\mathbf{x}}_{1:t})}\Big{[}D_{KL}(q({\mathbf{z}}^{s}_{t:1+T}|{\mathbf{z}}^{s}_{1:t})||p({\mathbf{z}}^{s}_{t+1:T}|{\mathbf{z}}^{s}_{1:t}))\Big{]}}_{\mathcal{L}_{KLD}^{s}}\\ &-\underbrace{D_{KL}(q({\mathbf{z}}^{e}_{1:t}|{\mathbf{x}}_{1:t})||p({\mathbf{z}}^{e}_{1:t}))-\mathbb{E}_{q({\mathbf{z}}^{e}_{1:t}|{\mathbf{x}}_{1:t})}\Big{[}D_{KL}(q({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{z}}^{e}_{1:t})||p({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{z}}^{e}_{1:t}))\Big{]}}_{\mathcal{L}_{KLD}^{e}}\end{split}$ (24)
@@ E.1 Identification of Latent Domain Variables ${\mathbf{u}}_{t}$
(Theorem 9 in (Allman et al.,, 2009)) Let $\mathbb{P}$ be a mixture in the form of Equation (26), such that for every $j$, the measures $\mu_{i,j}$ are linearly independent. Then, if $c\geq 3$, $\{\pi_{i},\mu_{i,j}\}$ are identifiable from $\mathbb{P}$ up to label swapping.
@@
The proof of this lemma can refer to Theorem 9 of (Allman et al.,, 2009). In general, Lemma E.1 shows that if the joint distribution of observation $\mathbb{P}$ can be decomposed into three linearly independent measures w.r.t. $\mu_{i,j}$ as shown in Equation (26), then the distributions of discrete latent variables are identifiable. Based on this Lemma, we further show the identification results of latent environments as follows.
@@
(Block-wise identifiability of the nonstationary latent variables ${\mathbf{z}}_{t}^{e}$ and the stationary latent variables ${\mathbf{z}}_{t}^{s}$. ) We follow the data generation process in Figure 2 and Equation (1)-(3), then we make the following assumptions:
  • A1 (Smooth and Positive Density:) The probability density function of latent variables is smooth and positive, i.e., $p({\mathbf{z}}^{e}_{t}|{\mathbf{z}}_{t-1}^{e},{\mathbf{z}}_{t-1}^{e})>0$ over $\mathcal{Z}_{t}^{e},\mathcal{Z}_{t-1}^{e}$ and $\mathcal{Z}_{t-2}^{e}$.
  • A2 (Linear Independent:) For any ${\mathbf{z}}_{t}^{e}\in\mathcal{Z}_{t}^{e}\subseteq\mathbb{R}^{n_{e}}$, ${\mathbf{v}}_{t-1,1},\cdots,{\mathbf{v}}_{t-1,n_{e}}$ as $n_{e}$ vector functions in $z_{t-2,1},\cdots,{\mathbf{v}}_{t-2,l},\cdots,z_{t-2,n_{e}}$ are linear independent, where ${\mathbf{v}}_{t-2,l}$ are formalized as follows:
    ${\mathbf{v}}_{t-2,l}=\frac{\partial\log p({\mathbf{z}}_{t}^{e}|{\mathbf{z}}_{t-1}^{e},{\mathbf{z}}_{t-2}^{e})}{\partial z_{t,k}^{e}\partial z_{t-2,l}^{s}}$ (27)
  • A3 (Domain Variability:) There exist two values of ${\mathbf{u}}=\{{\mathbf{z}}_{t-1}^{e},{\mathbf{z}}_{t-2}^{e}\}$, i.e., ${\mathbf{u}}_{1}$ and ${\mathbf{u}}_{2}$, s.t., for any set $\mathcal{A}_{{\mathbf{z}}_{t}}\subseteq\mathcal{Z}_{t}$ with non-zero probability measure and cannot be expressed as $B_{{\mathbf{z}}_{t}^{s}}\times{\mathbf{z}}_{t}^{e}$, for any $B_{{\mathbf{z}}_{t}^{s}}\subset\mathcal{Z}_{t}^{s}$, we have:
[…]
Then, by learning the data generation process, ${\mathbf{z}}_{t}^{e}$ are subspace identifiable.
Proof.
We start from the matched marginal distribution to develop the relation between ${\mathbf{z}}_{t}$ and $\hat{{\mathbf{z}}}_{t}$ as follows
For any ${\mathbf{z}}_{t-1}$ and ${\mathbf{z}}_{t-2}$, the Jacobian matrix of the mapping from $({\mathbf{x}}_{t-1},\hat{{\mathbf{z}}}_{t})$ to $({\mathbf{x}}_{t-1},{\mathbf{z}}_{t})$ is
$\begin{bmatrix}\mathbf{I}&\mathbf{0}\\ *&\mathbf{J}_{h}\\ \end{bmatrix},$
where * denotes a matrix, and the determinant of this Jacobian matrix is $|\mathbf{J}_{h}|$. Since ${\mathbf{x}}_{t-1}$ do not contain any information of $\hat{{\mathbf{z}}}_{t}$, the right-top element is $\mathbf{0}$. Therefore, $p(\hat{{\mathbf{z}}}_{t},{\mathbf{x}}_{t-1}|{\mathbf{x}}_{t-2})=p({\mathbf{z}}_{t},{\mathbf{x}}_{t-1}|{\mathbf{x}}_{t-2})\cdot|\mathbf{J}_{h}|$. Dividing both sides of this equation by $p({\mathbf{x}}_{t-1}|{\mathbf{e}}_{t})$ gives
$p(\hat{{\mathbf{z}}}_{t}|{\mathbf{x}}_{t-1},{\mathbf{x}}_{t-2})=p({\mathbf{z}}_{t}|{\mathbf{x}}_{t-1},{\mathbf{x}}_{t-2})\cdot|\mathbf{J}_{h}|.$ (30)
Since $p({{\mathbf{z}}}_{t}|{\mathbf{x}}_{t-1},{\mathbf{x}}_{t-2})=p({{\mathbf{z}}}_{t}|g({\mathbf{z}}_{t-1}),g({\mathbf{z}}_{t-2}))=p({{\mathbf{z}}}_{t}|{\mathbf{z}}_{t-1},{\mathbf{z}}_{t-2})$ and similarly $p(\hat{{\mathbf{z}}}_{t}|{\mathbf{x}}_{t-1},{\mathbf{x}}_{t-2})=p(\hat{{\mathbf{z}}}_{t}|{\mathbf{z}}_{t-1},{\mathbf{x}}_{t-2})$, we have:
$\begin{split}\log p(\hat{{\mathbf{z}}}_{t}|\hat{{\mathbf{z}}}_{t-1},{\mathbf{z}}_{t-2})&=\log p({\mathbf{z}}_{t}|{\mathbf{z}}_{t-1},{\mathbf{z}}_{t-2})+\log|\mathbf{J}_{h}|=\log p({\mathbf{z}}_{t}^{e}|{\mathbf{z}}_{t-1}^{e},{\mathbf{z}}_{t-2}^{e})+\log p({\mathbf{z}}_{t}^{s}|{\mathbf{z}}_{t-1}^{s})+\log|\mathbf{J}_{h}|.\end{split}$ (31)
Therefore, for $i\in\{n_{e}+1,\cdots,n\}$, the partial derivative of Equation (31) w.r.t $\hat{z}_{t,i}$ is
$\begin{split}\frac{\partial\log p(\hat{{\mathbf{z}}}_{t}|{{\mathbf{z}}}_{t-1},{{\mathbf{z}}}_{t-2})}{\partial\hat{z}_{t,i}}&=\frac{\partial\log p(\hat{{\mathbf{z}}}_{t}^{e}|{{\mathbf{z}}}_{t-1}^{e},{{\mathbf{z}}}_{t-2}^{e})}{\partial\hat{z}_{t,i}}+\frac{\partial\log p(\hat{{\mathbf{z}}}_{t}^{s}|{{\mathbf{z}}}_{t-1}^{s})}{\partial\hat{z}_{t,i}}\\ &=\sum_{k=1}^{n_{e}}\frac{\partial\log p({\mathbf{z}}_{t}^{e}|{\mathbf{z}}_{t-1}^{e},{\mathbf{z}}_{t-2}^{e})}{\partial z_{t,k}^{e}}\cdot\frac{\partial z_{t,k}^{e}}{\partial\hat{z}_{t,i}}+\sum_{k=n_{e}+1}^{n}\frac{\partial\log p(z^{s}_{t,k}|{\mathbf{z}}_{t-1}^{s})}{\partial z_{t,k}^{s}}\cdot\frac{\partial z_{t,k}^{s}}{\partial\hat{z}_{t,i}}+\frac{\partial\log|\mathbf{J_{h}}|}{\partial\hat{z}_{t,i}}.\end{split}$ (32)
Sequentially, for each $l=1,\cdots,n_{e}$, and each value of $z_{t-2,l}^{e}$, its partial derivative w.r.t. $z_{t-2,l}^{e}$ is shown as follows:
$\begin{split}&\frac{\partial\log p(\hat{{\mathbf{z}}}_{t}|{{\mathbf{z}}}_{t-1},{{\mathbf{z}}}_{t-2})}{\partial\hat{z}_{t,i}\partial z_{t-2,l}^{s}}=\frac{\partial\log p(\hat{{\mathbf{z}}}_{t}^{e}|{{\mathbf{z}}}_{t-1}^{e},{{\mathbf{z}}}_{t-2}^{e})}{\partial\hat{z}_{t,i}\partial z_{t-2,l}^{s}}+\frac{\partial\log p(\hat{{\mathbf{z}}}_{t}^{s}|{{\mathbf{z}}}_{t-1}^{s})}{\partial\hat{z}_{t,i}\partial z_{t-2,l}^{s}}\\ =&\sum_{k=1}^{n_{e}}\frac{\partial\log p({\mathbf{z}}_{t}^{e}|{\mathbf{z}}_{t-1}^{e},{\mathbf{z}}_{t-2}^{e})}{\partial z_{t,k}^{e}\partial z_{t-2,l}^{s}}\cdot\frac{\partial z_{t,k}^{e}}{\partial\hat{z}_{t,i}}+\sum_{k=n_{e}+1}^{n}\frac{\partial\log p(z^{s}_{t,k}|{\mathbf{z}}_{t-1}^{s})}{\partial z_{t,k}^{s}\partial z_{t-2,l}^{s}}\cdot\frac{\partial z_{t,k}^{s}}{\partial\hat{z}_{t,i}}+\frac{\partial\log|\mathbf{J_{h}}|}{\partial\hat{z}_{t,i}\partial z_{t-2,l}^{s}}.\end{split}$ (33)
Since the distribution $p(\hat{{\mathbf{z}}}_{t}^{e}|{\mathbf{z}}_{t-1}^{e},{\mathbf{z}}_{t-2}^{e})$ does not change across $\hat{z}_{t,i},i\in\{n_{e}+1,\cdots,n\}$, $\frac{\partial\log p(\hat{{\mathbf{z}}}_{t}^{e}|{\mathbf{z}}_{t-1}^{e},{\mathbf{z}}_{t-2}^{e})}{\partial\hat{z}_{t,i}}=0$. Since the distribution $p(\hat{{\mathbf{z}}}_{t}^{s}|{\mathbf{z}}_{t-1}^{s})$ does not change across different value of $z_{t-2,l}^{e}$, $\frac{\partial\log p(\hat{{\mathbf{z}}}_{t}^{s}|{{\mathbf{z}}}_{t-1}^{s})}{\partial\hat{z}_{t,i}\partial z_{t-2,l}}=0$. And given ${\mathbf{z}}_{t-1}^{s}$, ${\mathbf{z}}_{t-2}$ is independent of ${\mathbf{z}}_{t}^{s}$, so $\frac{\partial\log p(z_{t,k}^{s}|{\mathbf{z}}_{t-1}^{s})}{\partial z_{t,k}^{s}\partial z_{t-2,l}^{s}}=0$. Moreover, $\frac{\partial\log|\mathbf{J_{h}}|}{\partial\hat{z}_{t,i}\partial z_{t-2,l}^{s}}=0$, then Equation (33) can be rewritten as:
$0=\sum_{k=1}^{n_{e}}\frac{\partial\log p({\mathbf{z}}_{t}^{e}|{\mathbf{z}}_{t-1}^{e},{\mathbf{z}}_{t-2}^{e})}{\partial z_{t,k}^{e}\partial z_{t-2,l}^{s}}\cdot\frac{\partial z_{t,k}^{e}}{\partial\hat{z}_{t,i}}$ (34)
Note that $\frac{\partial z_{t,k}^{e}}{\partial\hat{z}_{t,i}}=0$ for $i=\{n_{e}+1,\cdots,n\}$ and $k\in\{1,\cdots,n_{d}\}$ means $\mathbf{B}=0$. Since $h(\cdot)$ is invertible, $\mathbf{J}_{h}$ is a full-rank matrix. Therefore, $\textbf{A}\neq 0$.
Besides, based on A3, one can show that all entries in the submatrix C zero according to part of the proof of Theorem 4.2 in (Kong et al.,, 2022)(Steps 1, 2, and 3). Therefore, ${\mathbf{z}}_{t}^{s}$ and ${\mathbf{z}}_{t}^{e}$ are block-wise identifiable. ∎
Theorem E.3.
(Identifiability of the latent environment $e_{t}.$) Suppose the observed data is generated following the data generation process in Figure 3 and Equation (1)-(3). Then we further make the following assumptions:
@@
  • A4 (Prior Environment Number:) The number of latent environments of the Markov process, $E$, is known.
  • A5 (Full Rank:) The transition matrix $\mathbf{A}$ is full rank.
@@
  • A6 (Linear Independence:) For $e=1,2,\cdots,E$, the probability measures $\mu_{e}=p({\mathbf{z}}_{t}^{e}|e_{t})$ are linearly independence and for any two different probability measures $\mu_{i},\mu_{j}$, their ratio $\frac{\mu_{i}}{\mu_{j}}$ are linearly independence.